# AllMatch: Exploiting All Unlabeled Data for Semi-Supervised Learning

### v1

@@ 1 Introduction
To address the trade-off between the quality and quantity of pseudo-labels in threshold-based pseudo-labeling, previous studies introduce dynamic threshold strategies that align with the evolving learning status of the model. For example, FlexMatch Zhang et al. (2021) utilizes the number of confident pseudo-labels to estimate the learning difficulty for each class, subsequently mapping the predefined threshold to class-specific thresholds based on the determined difficulty levels. Additionally, FreeMatch Wang et al. (2022) leverages the average confidence of unlabeled data to establish a dynamic global threshold. Besides, SoftMatch Chen et al. (2023) employs a Gaussian function representing the disparity between sample-specific and global confidence to model sample weights, assigning a positive weight to each unlabeled sample. However, samples with confidence significantly lower than the global confidence receive close-to-zero weights, essentially treated as if they were discarded. This characteristic positions SoftMatch as a variant of the threshold-based method. While the aforementioned algorithms effectively employ pseudo-labels to assess the learning status, the impact of biased data sampling or potential inter-class similarities can significantly influence the predictions of unlabeled data. Consequently, a pivotal question arises: Can we incorporate additional evidence along with pseudo-labels to achieve a more accurate estimation of the learning status?
@@ 2 Related Work
Consistency regulation represents the fundamental approach in SSL. The primary objective is to ensure consistent predictions across different perturbed views of the same sample. The original approaches Laine and Aila (2016); Sajjadi et al. (2016) minimize the L2 distance between the predictions of two randomly augmented views. The Mean Teacher method Tarvainen and Valpola (2017) extends this approach by replacing one of the predictions with the output of an exponential moving average (EMA) model. Furthermore, FixMatch Sohn et al. (2020) combines consistency regulation and pseudo-labeling techniques, establishing an effective SSL paradigm. To avoid similar perturbed views, Sohn et al. (2020); Xie et al. (2020) use the predictions of weakly augmented samples as pseudo-labels and minimize their divergence from the predictions of the corresponding strongly augmented views. Additionally, these methods use cross-entropy loss instead of the L2 distance for stable training.
@@
To ensure high-quality pseudo-labels, FixMatch Sohn et al. (2020) utilizes a high constant threshold throughout training to filter out potentially incorrect pseudo-labels. However, this strategy results in the underutilization of unlabeled data. To address this issue, FlexMatch Zhang et al. (2021) draws inspiration from curriculum learning Bengio et al. (2009), mapping the predefined threshold to class-specific thresholds based on the learning status of each category. Dash Xu et al. (2021) defines the threshold based on the loss of labeled data, eliminating the empirical threshold parameter. FreeMatch Wang et al. (2022) employs the average confidence on unlabeled data as the adaptive global threshold. SoftMatch Chen et al. (2023) estimates sample weights by a dynamic Gaussian function, maintaining soft margins between unlabeled samples of different confidence levels. Notably, samples with confidence significantly lower than the global confidence are assigned close-to-zero weights in SoftMatch. Accordingly, such samples can be considered dropped, rendering SoftMatch a variant of the threshold-based method. In addition to the threshold-based algorithms, CoMatch Li et al. (2021) and SimMatch Zheng et al. (2022) leverage contrastive loss to impose sample-level constraints on all unlabeled data. In contrast, AllMatch combines the advantages of both threshold-based and contrastive-based methods. The approach introduces CAT, a learning-status-aware threshold strategy, and BCC regulation, the semantic-level supervision for the entire unlabeled set, to maximize the utilization of unlabeled data and enhance the overall performance.
@@
In addition to consistency regulation and pseudo-labeling, entropy-based regulation is another widely adopted strategy. Entropy minimization Grandvalet and Bengio (2004) promotes high-confidence predictions during training. Maximizing the entropy of the expectation over all samples Krause et al. (2010); Arazo et al. (2020); Zhao et al. (2022) introduces the concept of fairness, which encourages the model to predict each class with equal frequency. Specifically, distribution alignment (DA) Berthelot et al. (2019a) and uniform alignment (UA) Chen et al. (2023) are prevailing strategies for achieving fairness in SSL, which adjusts the pseudo-label based on the overall predictions on the unlabeled data.
@@ 4.1 Balanced Semi-Supervised Learning
Performance. Table 1 presents the top-1 accuracy on CIFAR-10/100, SVHN, and STL-10 with various numbers of labeled samples. The performance on ImageNet is reported in Table 2. The experimental results demonstrate that AllMatch achieves state-of-the-art performance on all datasets, except for CIFAR-100 with 10000 labels and SVHN with 40 labels, where our model’s performance remains comparable to the best competitor. Moreover, regarding CIFAR-100, AllMatch outperforms ReMixMatch when only 400 or 2500 labels are available, while the latter achieves better performance when 10000 labels are available. The competitive results obtained by ReMixMatch mainly stem from the Mixup technique Zhang et al. (2017) and the additional self-supervised learning part. Furthermore, AllMatch exhibits substantial advantages over previous algorithms when dealing with extremely limited labeled samples. Specifically, the approach surpasses the second-best counterpart by 1.87% on CIFAR-10 with 10 labels, 0.66% on CIFAR-100 with 400 labels, and 2.86% on STL-10 with 40 labels. Particularly, STL-10 poses significant challenges due to its large unlabeled set that comprises 100k images. Accordingly, the impressive improvement obtained on STL-10 highlights the potential of AllMatch to be deployed in real-world applications.
@@ 4.2 Imbalanced Semi-Supervised Learning
Settings. We evaluate AllMatch in the context of imbalanced SSL, where both labeled and unlabeled data exhibit a long-tailed distribution. All experiments are conducted on the TorchSSL codebase. Following prior studies Lee et al. (2021); Oh et al. (2022); Wei et al. (2021); Lai et al. (2022); Fan et al. (2022); Chen et al. (2023), we generate the labeled and unlabeled sets using the configurations of $N_{c}=N_{1}\cdot\gamma^{-\frac{c-1}{C-1}}$ and $M_{c}=M_{1}\cdot\gamma^{-\frac{c-1}{C-1}}$. Specifically, for CIFAR-10-LT, we set $N_{1}$ to $1500$, $M_{1}$ to $3000$, and $\gamma$ to range from 50 to 150. For CIFAR-100-LT, we set $N_{1}$ to $150$, $M_{1}$ to $300$, and $\gamma$ to range from 20 to 100. In all experiments, we employ WRN-28-2 as the backbone and utilize the Adam optimizer Kingma and Ba (2014) with the weight decay of 4e-5. The batch sizes $B_{L}$ and $B_{U}$ are set to 64 and 128, respectively. The learning rate is initially set to 2e-3 and adjusted by a cosine decay scheduler during training. We repeat each experiment three times and report the overall performance. Detailed implementation is listed in Appendix C.

### v2

@@ 1 Introduction
To address the trade-off between the quality and quantity of pseudo-labels in threshold-based pseudo-labeling, previous studies introduce dynamic threshold strategies that align with the evolving learning status of the model. For example, FlexMatch Zhang et al. (2021) utilizes the number of confident pseudo-labels to estimate the learning difficulty for each class, subsequently mapping the predefined threshold to class-specific thresholds based on the determined difficulty levels. Additionally, FreeMatch Wang et al. (2022) leverages the average confidence of unlabeled data to establish a dynamic global threshold. Besides, SoftMatch Chen et al. (2023a) employs a Gaussian function representing the disparity between sample-specific and global confidence to model sample weights, assigning a positive weight to each unlabeled sample. However, samples with confidence significantly lower than the global confidence receive close-to-zero weights, essentially treated as if they were discarded. This characteristic positions SoftMatch as a variant of the threshold-based method. While the aforementioned algorithms effectively employ pseudo-labels to assess the learning status, the impact of biased data sampling or potential inter-class similarities can significantly influence the predictions of unlabeled data. Consequently, a pivotal question arises: Can we incorporate additional evidence along with pseudo-labels to achieve a more accurate estimation of the learning status?
@@ 2 Related Work
Consistency regulation and pseudo labelling Laine and Aila (2016); Sajjadi et al. (2016); Tarvainen and Valpola (2017) are fundamental approaches in SSL. The former encourages consistent predictions across different perturbed views of the same sample while the latter assigns pseudo-labels to unlabeled samples. Among established techniques, FixMatch Sohn et al. (2020) perfectly combines both techniques, establishing an effective SSL paradigm. Specifically, FixMatch uses the predictions of weakly augmented samples as pseudo-labels and minimize their divergence from the predictions of the corresponding strongly augmented views.
@@
To ensure high-quality pseudo-labels, FixMatch utilizes a high constant threshold throughout training to filter out potentially incorrect pseudo-labels. However, this strategy results in the underutilization of unlabeled data. To address this issue, FlexMatch Zhang et al. (2021) draws inspiration from curriculum learning Bengio et al. (2009), mapping the predefined threshold to class-specific thresholds based on the learning status of each category. Dash Xu et al. (2021) defines the threshold based on the loss of labeled data, eliminating the empirical threshold parameter. FreeMatch Wang et al. (2022) employs the average confidence on unlabeled data as the adaptive global threshold. SoftMatch Chen et al. (2023a) estimates sample weights by a dynamic Gaussian function, maintaining soft margins between unlabeled samples of different confidence levels. In addition to the threshold-based algorithms, CoMatch Li et al. (2021) and SimMatch Zheng et al. (2022) leverage contrastive loss to impose sample-level constraints on all unlabeled data. In contrast, AllMatch combines the advantages of both threshold-based and contrastive-based methods by introducing CAT, a learning-status-aware threshold strategy, and BCC regulation, the semantic-level supervision for the entire unlabeled set.
Concurrent with our work, FullMatch Chen et al. (2023b) also introduces semantic guidance for all unlabeled samples. Specifically, FullMatch compares the predictions of weakly and strongly augmented samples to identify negative classes. In contrast, AllMatch identifies candidate classes by comparing sample and global top-k confidence, thus considering the learning state of both individual sample and model in this process. Additionally, FullMatch assigns low probability (similar to label smoothing) for the negative classes in the optimization objective, while AllMatch directly encourages consistent candidate-negative division for all unlabeled samples, thus exhibiting better consistency with the unsupervised loss.
@@
In addition to consistency regulation and pseudo-labeling, entropy-based regulation is another widely adopted strategy. Entropy minimization Grandvalet and Bengio (2004) promotes high-confidence predictions during training. Maximizing the entropy of the expectation over all samples Krause et al. (2010); Arazo et al. (2020); Zhao et al. (2022) introduces the concept of fairness, which encourages the model to predict each class with equal frequency. Specifically, distribution alignment (DA) Berthelot et al. (2019a) and uniform alignment (UA) Chen et al. (2023a) are prevailing strategies for achieving fairness in SSL, which adjusts the pseudo-label based on the overall predictions on the unlabeled data.
@@ 4.1 Balanced Semi-Supervised Learning
Performance. Table 1 presents the top-1 accuracy on CIFAR-10/100, SVHN, and STL-10 with various numbers of labeled samples. The performance on ImageNet is reported in Table 2. The experimental results demonstrate that AllMatch achieves state-of-the-art performance on almost all datasets. For CIFAR-10, AllMatch outperforms FullMatch with only 40 available labels, and performs comparably to FullMatch when there are 250 or 4000 labels. Moreover, regarding CIFAR-100, AllMatch outperforms ReMixMatch when only 400 or 2500 labels are available, while the latter achieves better performance when 10000 labels are available. The competitive results obtained by ReMixMatch mainly stem from the Mixup technique Zhang et al. (2017) and the additional self-supervised learning part. Furthermore, AllMatch exhibits substantial advantages over previous algorithms when dealing with extremely limited labeled samples. Specifically, the approach surpasses the second-best counterpart by 1.87% on CIFAR-10 with 10 labels, 0.66% on CIFAR-100 with 400 labels, and 2.86% on STL-10 with 40 labels. Particularly, STL-10 poses significant challenges due to its large unlabeled set that comprises 100k images. Accordingly, the impressive improvement obtained on STL-10 highlights the potential of AllMatch to be deployed in real-world applications.
@@ 4.2 Imbalanced Semi-Supervised Learning
Settings. We evaluate AllMatch in the context of imbalanced SSL, where both labeled and unlabeled data exhibit a long-tailed distribution. All experiments are conducted on the TorchSSL codebase. Following prior studies Lee et al. (2021); Oh et al. (2022); Wei et al. (2021); Lai et al. (2022); Fan et al. (2022); Chen et al. (2023a), we generate the labeled and unlabeled sets using the configurations of $N_{c}=N_{1}\cdot\gamma^{-\frac{c-1}{C-1}}$ and $M_{c}=M_{1}\cdot\gamma^{-\frac{c-1}{C-1}}$. Specifically, for CIFAR-10-LT, we set $N_{1}$ to $1500$, $M_{1}$ to $3000$, and $\gamma$ to range from 50 to 150. For CIFAR-100-LT, we set $N_{1}$ to $150$, $M_{1}$ to $300$, and $\gamma$ to range from 20 to 100. In all experiments, we employ WRN-28-2 as the backbone and utilize the Adam optimizer Kingma and Ba (2014) with the weight decay of 4e-5. The batch sizes $B_{L}$ and $B_{U}$ are set to 64 and 128, respectively. The learning rate is initially set to 2e-3 and adjusted by a cosine decay scheduler during training. We repeat each experiment three times and report the overall performance. Detailed implementation is listed in Appendix C.